OpenStudy (anonymous):

Solve using the square root method. Round the solution to the nearest tenth. x2 - 20 = 0

OpenStudy (anonymous):

\[x^2 - 20 = 0\]\[x^2 = 20\]\[x = \pm \sqrt{20}\]